A rectangular paper has width of 8in and length 11in.A typist leaves a one inch margin on both sides of the width and 1.5in margin on both side of length and uses remaining middle part for typing.What fractional part of the page's area is being used for typing?
Answer by Theo(13342)

the paper has a width of 8 inches and a length of 11 inches.

the total area is therefore equal to 8 * 11 = 88 square inches.

the typist leaves 1 inch on both sides of the width and 1.5 inches on both sides of the length.

the width is therefore reduced by 2 inches and the length is therefore reduced by 3 inches.

the width for typing is equal ti 8 - 2 = 6 inches.
the length for typing is equal to 11 - 3 = 8 inches.

the area for typins is therefore equal to 6 * 8 = 48 square inches.

the fractional part of the page's area that is being used for typing is therefore equal to 48 / 88 = 24 / 44 = 12 / 22 = 6 / 11.

the decimal equivalent of 6 / 11 is .54..... with the 54 repeated endlessly.